An axe that weighs 8.0 kg is used to strikes the log with a force of 400 N that comes to rest in 0.04 seconds. What is the magnitude of the impulse? How fast is the axe approach the log?
1
Expert's answer
2020-11-05T10:45:04-0500

By definition, the impulse is:


"J = F\Delta t"

where "F = 400N" is the force and "\Delta t = 0.04s" is the time of interaction. Thus:


"J = 400N\times 0.04s = 16\space kg\cdot m/s"

According to the second Newton's law, the acceleration of the axe is:


"a = \dfrac{F}{m} = \dfrac{400N}{8kg} = 50m/s^2"

where "m = 8kg" is the mass of the axe.

On the other hand, by definition, the acceleration is:


"a = \left| \dfrac{v_f - v_i}{\Delta t} \right|"

where "v_f" is the final speed of the axe and "v_i" is the initial speed of the axe. Since the axe eventually comes to rest, its final speed is "v_f = 0". Then the initial speed is:


"v_i = a\Delta t = 50m/s^2\times 0.04s = 2m/s"

Answer. "J = 16 \space kg\cdot m/s", "v_i = 2m\/s".